OpenStudy (anonymous):

The hypotenuse of a right triangle is 26 mm. One leg of the right triangle is 10 mm. What is the length of the other leg?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

plzzzzzzzzzzzzzz help

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (anonymous):

sill fan and meadal

OpenStudy (arindameducationusc):

Use pythagoras theorem..... Pythagoras' theorem, is a relation in Euclidean geometry among the three sides of a right triangle. It states that the square of the hypotenuse (the side opposite the right angle) is equal to the sum of the squares of the other two sides.

OpenStudy (arindameducationusc):

26^2=10^2+x^2 Find x after that. :)

OpenStudy (arindameducationusc):

See this

OpenStudy (arindameducationusc):

Got it?

OpenStudy (anonymous):

a little

OpenStudy (sepeario):

Basically the concept is if you have three sides in a right-angled triangle, you can assume that the first side squared plus the second side squared equals the third side squared.

OpenStudy (arindameducationusc):

@Sepeario will explain you dont worry......

OpenStudy (sepeario):

Meaning that in the triangle that you are given, \[10^2+x^2=26^2\] where x is the side we are trying to find the length of.

OpenStudy (sepeario):

Can you try to solve the equation I have given above for x?

OpenStudy (anonymous):

kk

OpenStudy (sepeario):

Have you solved it yet?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

i am very short on time

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (anonymous):

are u there ??????

OpenStudy (sepeario):

What don't you understand?

OpenStudy (anonymous):

the answer

OpenStudy (sepeario):

Can you solve the equation \[10^2+x^2=26^2\]

OpenStudy (sepeario):

That's a very simple algebraical equation which you should be capable of evaluating at your level I think.